OpenStudy (anonymous):

(f+g)(x) (fxg)(x) (f-g)(2) f(x)=2x squared g(X)=3x

OpenStudy (imtiaz7):

Question???

OpenStudy (anonymous):

write the expression for the first 3 and evaluate .. the functions for the 2 numbers under them ... thank you so much for your help

OpenStudy (imtiaz7):

(f+g)(x)=f(x)+g(x)=2x^2+3x follow this pattern and try to solve others, otherwise i'll solve these lol